OpenStudy (anonymous):

6x + 5y = 10 7x – 3y = 24 ?

OpenStudy (anonymous):

Solve each system of equations

OpenStudy (cydney_morgan):

6x + 5y = 10 7x – 3y = 24 ----- Multiply top equation by 3 Multiply bottom equation by 5 ------ 18x + 15y = 30 35x - 15y = 120 ---- Add and solve for "x") 53x = 150 x = 150/53 ----------------- Solve for "y" using 6x + 5y = 10 ----- 6(150/53) + 5y = 10 900/53 + 5y = 10 ---- 900/53 + 5y = 530/53 ------ 5y = -370/53 y = -370/265

OpenStudy (anonymous):

(2.83, -1.4)